# Ground state solutions to the nonlinear Born-Infeld problem

### v1

@@ 1. Introduction
and $u$ is a weak solution of (1.1) provided $u$ is a critical point of $I$ with $\|\nabla u\|_{\infty}<1$. However, it should be pointed out that for the well-definedness of $\Psi$, the condition $|\nabla u|\leq 1$ for a.e. $x\in\mathbb{R}^{N}$ is required and it is not possible to find a function space on which $\Psi$ becomes of class ${\mathcal{C}}^{1}$. Therefore, we need to go beyond the smooth critical point theory to treat $I$ directly on some function space, and the nonsmoothness of $\Psi$ makes the problem complicated.
@@
1.1. The zero mass case
@@
By (1.4), (F1) and (F2), we may verify $\Phi\in{\mathcal{C}}^{1}(X_{0};\mathbb{R})$ and it allows us to define the notion of a critical point of $I$ as in [32]. Namely, $u\in X_{0}$ is said to be a critical point of $I$ if
@@ 1.2. Positive mass case
Then, the variational functional $I$ is given by (1.2), $\Psi$ is defined as before with the natural domain
@@
Remark that $\Psi$ is a convex functional on a convex domain $D(\Psi)$ and $\Phi\in{\mathcal{C}}^{1}(X_{1};\mathbb{R})$. Therefore, the notation of critical points of $I$ on $X_{1}$ is defined as in (1.5) by changing $X_{0}$ to $X_{1}$. We also introduce the Pohožaev set as
@@ 1.4. Comments
Even though we know the existence of minimizers, it is worth noting that whether these minimizers become critical points of $I$ is not trivial or immediate, and this point is more delicate than the semilinear case. For the semilinear case, a similar assertion is proved in [29, Lemma 3.3] where the differentiability of functional originated from the differential operator is utilized. In our case, the corresponding functional is $\Psi$, which is not smooth on $X_{0}$ or $X_{1}$. In addition, the nonsmoothness of $M_{0}$ and $M_{1}$ affect the properties of the projections to ${\mathcal{M}}_{0}$ and ${\mathcal{M}}_{1}$, and the analysis of the projections onto ${\mathcal{M}}_{0}$, ${\mathcal{M}}_{1}$ and functional $I$ is more delicate and complicated. These issues will be solved in Subsections˜4.3 and 5.2 and it is one of novelties in this paper.
@@
It is also worth mentioning that our argument may be extended to a more general problem, for instance $f(x,u)$ in place of $f(u)$ in (1.1). In fact, for the semilinear case, the characterization of the ground state energy by the mountain pass level is obtained in [23], and it is useful to handle the nonlinearity $f(x,u)$ or a potential term $V(x)u$. The same characterization for (1.1) is obtained in Theorems˜1.2 and 1.6, and we expect that this might be helpful to study (1.1) with $f(x,u)$ or $V(x)u$. Furthermore, though it is only for minimizers, we succeed to prove that the constraint $M_{0}(u)=0$ (resp. $M_{1}(u)=0$) is natural in the sense that a critical point of $I|_{{\mathcal{M}}_{0}}$ (resp. $I|_{{\mathcal{M}}_{1}}$) provides a critical point of $I$. Therefore, it is interesting to see whether the same is true for higher energies or other nonsmooth constraints.
The structure of the paper is as follows. In Section˜2, the functional setting is recalled and some preliminary properties related to the functionals $I$ and $\Psi$ are proved. In Section˜3 we develop profile decompositions for bounded sequences in $X_{0}$ and $X_{1}$. Section˜4 is then devoted to the zero mass case and contains the proofs of Theorems˜1.2 and 1.4. In Section˜5 we turn to the positive mass case and prove Theorem˜1.6. Finally, in Section˜6 we establish the existence of nonradial solutions.
@@ 2. Concentration-compactness and a profile decomposition
Finally, the space dimension $N$ is assumed to be $N\geq 3$ in the zero mass case and $N\geq 1$ in the positive mass case.
@@
Finally, we prove that every critical point of $I$ in $X$ becomes a classical solution of (1.1) and satisfies the Pohožaev identity. A similar assertion is obtained in [18, Propositions 2.6 and 2.10] for $N\geq 3$. Here our main focus is in the case $N=1$. Notice that under (F1)–(F2) (resp. (G1)), $\Phi\in{\mathcal{C}}^{1}(X;\mathbb{R})$ due to (2.2).
@@
Suppose either $N\geq 3$ and (F1)–(F2) or else $N\geq 1$ and (G1). Then each critical point $u$ of $I$ in $X$ is a classical solution of (1.1) and the Pohožaev identity $M(u)=0$ holds.
@@
Let $u\in X$ be a critical point of $I$ and define
@@
Let $N=1$ and $u\in X$ be a critical point of $I$, and notice that $\rho\in L^{2}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$ in view of (G1) and (2.2). We follow the argument in [8, Theorem 1.4, Proposition 2.7, Remark 2.9]. Without any changes, the argument in [8, Proposition 2.7] shows that
@@ 3. Concentration-compactness and a profile decomposition
In this section, we establish a profile decomposition for bounded sequences in $X$, which is useful to prove the existence of minimizers of $I$ on ${\mathcal{M}}_{i}$. Recall that in this section, the same notation to Section˜2 is used: $p_{0}$, $X$ and $\Psi_{0}$. For the profile decomposition, several lemmas are necessary. Firstly, a slight extension of [18, Lemma 2.27] is needed:
@@ 4.3. Minimizers on the Pohožaev set are critical points
The aim of this subsection is to prove that every minimizer corresponding to $c_{0}$ is a critical point of $I$, and hence Section˜2 implies that it is a classical solution of (1.1).
@@
Let $u_{0}\in{\mathcal{M}}_{0}$ satisfy $I(u_{0})=c_{0}$. Then $u_{0}$ is a critical point of $I$.
@@
To prove that $u_{0}$ is a critical point of $I$, let $v\in D(\Psi)$ be any point: otherwise, (1.5) holds trivially. For $t\in(0,1)$, we introduce
@@
and recalling that $u_{0}$ minimizes $I$ on ${\mathcal{M}}_{0}$, we see from (4.8) that
@@
Thus (4.9) does not occur and $u_{0}$ is a critical point of $I$. ∎
4.4. Existence of a minimizer and other properties of minimizers
@@
Thus $\theta_{j_{0}}=1$, $\widetilde{u}_{j_{0}}\in{\mathcal{M}}_{0}$, $\lim_{n\to\infty}\int_{\mathbb{R}^{N}}f(u_{n})u_{n}\,dx=\int_{\mathbb{R}^{N}}f(\widetilde{u}_{j_{0}})\widetilde{u}_{j_{0}}\,dx$ and $I(\widetilde{u}_{j_{0}})=c_{0}$. Moreover, since $f(s)s>0$ for $s\neq 0$ in view of (F3) and (F5), by (4.12), $\widetilde{u}_{j}=0$ for each $j\neq j_{0}$, $J=\{j_{0}\}$ and $\lim_{n\to\infty}\int_{\mathbb{R}^{N}}F(u_{n})\,dx=\int_{\mathbb{R}^{N}}F(\widetilde{u}_{j_{0}})\,dx$. Furthermore, by $I(u_{n})\to c_{0}=I(\widetilde{u}_{j_{0}})$, we have $\Psi(u_{n})\to\Psi(\widetilde{u}_{j_{0}})$, and Section˜2 yields $\|u_{n}(\cdot+y_{n}^{j_{0}})-\widetilde{u}_{j_{0}}\|_{X_{0}}\to 0$. Thanks to Subsection˜4.3 we know that $\widetilde{u}_{j_{0}}$ is a critical point of $I$ and Section˜2 implies that $\widetilde{u}_{j_{0}}$ is a classical solution of (1.1).
@@ 5.2. Minimizers on the Pohožaev set are critical points
To prove that minimizers on ${\mathcal{M}}_{1}$ are critical points of $I$, we need the following:
@@
Lemma 5.5.
Let $u_{0}\in{\mathcal{M}}_{1}$ satisfy $I(u_{0})=c_{1}$. Then $u_{0}$ is a critical point of $I$.
@@
By $0<c_{1}=I(u_{0})=\int_{\mathbb{R}^{N}}f(u_{0})u_{0}\,dx/N$, we can repeat the same reasoning as for the zero mass case in Subsection˜4.3 to show that every minimizer of $I$ on ${\mathcal{M}}_{1}$ is a critical point. ∎
@@
is bounded and that each $u_{n}$ is $1$-Lipschitz in view of $u_{n}\in D(\Psi)$, we observe from (5.2) and (5.6) that $(u_{n})_{n\in\mathbb{N}}$ is bounded in $L^{\infty}(\mathbb{R}^{N})$. Furthermore, by (5.2) and (5.6), for each $t>0$ there exists $\delta_{t}>0$ such that $g(s)s-2G(s)\geq\delta_{t}$ if $|s|\geq t$. Therefore, $(|\{|u_{n}|\geq t\}|)_{n\in\mathbb{N}}$ is also bounded where $|A|$ denotes the $N$-dimensional Lebesgue measure of $A\subset\mathbb{R}^{N}$. By (G1), choose $t_{0}>0$ so that $|G(t)|\leq t^{2}/4$ for each $|t|\leq t_{0}$. Then
@@
Therefore, $\theta_{j_{0}}=1$ and $\widetilde{u}_{j_{0}}\in{\mathcal{M}}_{1}$ is a minimizer. Furthermore, by Subsection˜5.2, $\widetilde{u}_{j_{0}}$ is a critical point of $I$.
@@ 6. Nonradial solutions
where $X=X_{0}$ in the zero mass case and $X=X_{1}$ in the positive mass case. Since $f$ (resp. $g$) is odd, $\Psi(-u)=\Psi(u)$ and $\Phi(-u)=\Phi(u)$. According to the principle of symmetric criticality (see [27, 18]), any critical point of $I$ in $X_{\tau}$ is also a critical point in $X_{\mathcal{O}}$, and by applying the principle again, it is a critical point in $X$. Thus, it suffices to find a critical point of $I|_{X_{\tau}}$.
@@
Let $u\in{\mathcal{M}}_{\tau}$ satisfy $I(u)=\inf_{{\mathcal{M}}_{\tau}}I$. Then we may use the argument in the proof of Subsections˜4.3 and 5.2 to verify that $u$ is a critical point of $I$ on $X_{\tau}$. As pointed in the beginning of this section, $u$ is a critical point of $I$ in $X$, and hence Section˜2 implies that $u$ is a classical solution of (1.1).

### v2

@@ 1. Introduction
and $u$ is a weak solution of (1.1) provided $u$ is a critical point of II with $\|\nabla u\|_{\infty}<1$. However, it should be pointed out that for the well-definedness of $\Psi$, the condition $|\nabla u|\leq 1$ for a.e. $x\in\mathbb{R}^{N}$ is required and it is not possible to find a function space on which $\Psi$ becomes of class ${\mathcal{C}}^{1}$. Therefore, we need to go beyond the smooth critical point theory to treat II directly on some function space, and the nonsmoothness of $\Psi$ makes the problem complicated.
@@
$1.$ The zero mass case
@@
By (1.4), (F1) and (F2), we may verify $\Phi\in{\mathcal{C}}^{1}(X_{0};\mathbb{R})$ and it allows us to define the notion of a critical point of II as in [32]. Namely, $u\in X_{0}$ is said to be a critical point of II if
@@ 1.2. Positive mass case
Then, the variational functional II is given by (1.2), $\Psi$ is defined as before with the natural domain
@@
Remark that $\Psi$ is a convex functional on a convex domain $D(\Psi)$ and $\Phi\in{\mathcal{C}}^{1}(X_{1};\mathbb{R})$. Therefore, the notation of critical points of II on $X_{1}$ is defined as in (1.5) by changing $X_{0}$ to $X_{1}$. We also introduce the Pohožaev set as
@@ 1.4. Comments
Even though we know the existence of minimizers, it is worth noting that whether these minimizers become critical points of II is not trivial or immediate, and this point is more delicate than the semilinear case. For the semilinear case, a similar assertion is proved in [29, Lemma 3.3] where the differentiability of functional originated from the differential operator is utilized. In our case, the corresponding functional is $\Psi$, which is not smooth on $X_{0}$ or $X_{1}$. In addition, the nonsmoothness of $M_{0}$ and $M_{1}$ affect the properties of the projections to ${\mathcal{M}}_{0}$ and ${\mathcal{M}}_{1}$, and the analysis of the projections onto ${\mathcal{M}}_{0}$, ${\mathcal{M}}_{1}$ and functional II is more delicate and complicated. These issues will be solved in Subsections˜4.3 and 5.2 and it is one of novelties in this paper.
@@
It is also worth mentioning that our argument may be extended to a more general problem, for instance $f(x,u)$ in place of $f(u)$ in (1.1). In fact, for the semilinear case, the characterization of the ground state energy by the mountain pass level is obtained in [23], and it is useful to handle the nonlinearity $f(x,u)$ or a potential term $V(x)u$. The same characterization for (1.1) is obtained in Theorems˜1.2 and 1.6, and we expect that this might be helpful to study (1.1) with $f(x,u)$ or $V(x)u$. Furthermore, though it is only for minimizers, we succeed to prove that the constraint $M_{0}(u)=0$ (resp. $M_{1}(u)=0$) is natural in the sense that a critical point of $I|_{{\mathcal{M}}_{0}}$ (resp. $I|_{{\mathcal{M}}_{1}}$) provides a critical point of II. Therefore, it is interesting to see whether the same is true for higher energies or other nonsmooth constraints.
The structure of the paper is as follows. In Section˜2, the functional setting is recalled and some preliminary properties related to the functionals II and $\Psi$ are proved. In Section˜3 we develop profile decompositions for bounded sequences in $X_{0}$ and $X_{1}$. Section˜4 is then devoted to the zero mass case and contains the proofs of Theorems˜1.2 and 1.4. In Section˜5 we turn to the positive mass case and prove Theorem˜1.6. Finally, in Section˜6 we establish the existence of nonradial solutions.
@@ 2. Concentration-compactness and a profile decomposition
Finally, the space dimension NN is assumed to be $N\geq 3$ in the zero mass case and $N\geq 1$ in the positive mass case.
@@
Finally, we prove that every critical point of II in $X$ becomes a classical solution of (1.1) and satisfies the Pohožaev identity. A similar assertion is obtained in [18, Propositions 2.6 and 2.10] for $N\geq 3$. Here our main focus is in the case $N=1$. Notice that under (F1)–(F2) (resp. (G1)), $\Phi\in{\mathcal{C}}^{1}(X;\mathbb{R})$ due to (2.2).
@@
Suppose either $N\geq 3$ and (F1)–(F2) or else $N\geq 1$ and (G1). Then each critical point $u$ of II in $X$ is a classical solution of (1.1) and the Pohožaev identity $M(u)=0$ holds.
@@
Let $u\in X$ be a critical point of II and define
@@
Let $N=1$ and $u\in X$ be a critical point of II, and notice that $\rho\in L^{2}(\mathbb{R})\cap L^{\infty}(\mathbb{R})$ in view of (G1) and (2.2). We follow the argument in [8, Theorem 1.4, Proposition 2.7, Remark 2.9]. Without any changes, the argument in [8, Proposition 2.7] shows that
@@ 3. Concentration-compactness and a profile decomposition
In this section, we establish a profile decomposition for bounded sequences in $X$, which is useful to prove the existence of minimizers of II on ${\mathcal{M}}_{i}$. Recall that in this section, the same notation to Section˜2 is used: $p_{0}$, $X$ and $\Psi_{0}$. For the profile decomposition, several lemmas are necessary. Firstly, a slight extension of [18, Lemma 2.27] is needed:
@@ 4.3. Minimizers on the Pohožaev set are critical points
The aim of this subsection is to prove that every minimizer corresponding to $c_{0}$ is a critical point of II, and hence Section˜2 implies that it is a classical solution of (1.1).
@@
Let $u_{0}\in{\mathcal{M}}_{0}$ satisfy $I(u_{0})=c_{0}$. Then $u_{0}$ is a critical point of II.
@@
To prove that $u_{0}$ is a critical point of II, let $v\in D(\Psi)$ be any point: otherwise, (1.5) holds trivially. For $t\in(0,1)$, we introduce
@@
and recalling that $u_{0}$ minimizes II on ${\mathcal{M}}_{0}$, we see from (4.8) that
@@
Thus (4.9) does not occur and $u_{0}$ is a critical point of II. ∎
$4.$ Existence of a minimizer and other properties of minimizers
@@
Thus $\theta_{j_{0}}=1$, $\widetilde{u}_{j_{0}}\in{\mathcal{M}}_{0}$, $\lim_{n\to\infty}\int_{\mathbb{R}^{N}}f(u_{n})u_{n}\,dx=\int_{\mathbb{R}^{N}}f(\widetilde{u}_{j_{0}})\widetilde{u}_{j_{0}}\,dx$ and $I(\widetilde{u}_{j_{0}})=c_{0}$. Moreover, since $f(s)s>0$ for $s\neq 0$ in view of (F3) and (F5), by (4.12), $\widetilde{u}_{j}=0$ for each $j\neq j_{0}$, $J=\{j_{0}\}$ and $\lim_{n\to\infty}\int_{\mathbb{R}^{N}}F(u_{n})\,dx=\int_{\mathbb{R}^{N}}F(\widetilde{u}_{j_{0}})\,dx$. Furthermore, by $I(u_{n})\to c_{0}=I(\widetilde{u}_{j_{0}})$, we have $\Psi(u_{n})\to\Psi(\widetilde{u}_{j_{0}})$, and Section˜2 yields $\|u_{n}(\cdot+y_{n}^{j_{0}})-\widetilde{u}_{j_{0}}\|_{X_{0}}\to 0$. Thanks to Subsection˜4.3 we know that $\widetilde{u}_{j_{0}}$ is a critical point of II and Section˜2 implies that $\widetilde{u}_{j_{0}}$ is a classical solution of (1.1).
@@ 5.2. Minimizers on the Pohožaev set are critical points
To prove that minimizers on ${\mathcal{M}}_{1}$ are critical points of II, we need the following:
@@
Lemma $5.$
Let $u_{0}\in{\mathcal{M}}_{1}$ satisfy $I(u_{0})=c_{1}$. Then $u_{0}$ is a critical point of II.
@@
By $0<c_{1}=I(u_{0})=\int_{\mathbb{R}^{N}}f(u_{0})u_{0}\,dx/N$, we can repeat the same reasoning as for the zero mass case in Subsection˜4.3 to show that every minimizer of II on ${\mathcal{M}}_{1}$ is a critical point. ∎
@@
is bounded and that each $u_{n}$ is 11-Lipschitz in view of $u_{n}\in D(\Psi)$, we observe from (5.2) and (5.6) that $(u_{n})_{n\in\mathbb{N}}$ is bounded in $L^{\infty}(\mathbb{R}^{N})$. Furthermore, by (5.2) and (5.6), for each $t>0$ there exists $\delta_{t}>0$ such that $g(s)s-2G(s)\geq\delta_{t}$ if $|s|\geq t$. Therefore, $(|\{|u_{n}|\geq t\}|)_{n\in\mathbb{N}}$ is also bounded where $|A|$ denotes the NN-dimensional Lebesgue measure of $A\subset\mathbb{R}^{N}$. By (G1), choose $t_{0}>0$ so that $|G(t)|\leq t^{2}/4$ for each $|t|\leq t_{0}$. Then
@@
Therefore, $\theta_{j_{0}}=1$ and $\widetilde{u}_{j_{0}}\in{\mathcal{M}}_{1}$ is a minimizer. Furthermore, by Subsection˜5.2, $\widetilde{u}_{j_{0}}$ is a critical point of II.
@@ 6. Nonradial solutions
where $X=X_{0}$ in the zero mass case and $X=X_{1}$ in the positive mass case. Since $f$ (resp. $g$) is odd, $\Psi(-u)=\Psi(u)$ and $\Phi(-u)=\Phi(u)$. According to the principle of symmetric criticality (see [27, 18]), any critical point of II in $X_{\tau}$ is also a critical point in $X_{\mathcal{O}}$, and by applying the principle again, it is a critical point in $X$. Thus, it suffices to find a critical point of $I|_{X_{\tau}}$.
@@
Let $u\in{\mathcal{M}}_{\tau}$ satisfy $I(u)=\inf_{{\mathcal{M}}_{\tau}}I$. Then we may use the argument in the proof of Subsections˜4.3 and 5.2 to verify that $u$ is a critical point of II on $X_{\tau}$. As pointed in the beginning of this section, $u$ is a critical point of II in $X$, and hence Section˜2 implies that $u$ is a classical solution of (1.1).